The surface area in feet of a cylinder with a radius of 1 foot can be found by adding the product of 2π and the square of the radius to the product of 2π and the radius. How does the surface area change when the radius is tripled from 1 foot to 3 feet?
•how does the area of a square change when the side length is tripled?
•the surface area of a sphere is the product of 4π and the square of the radius. How does the surface area change when the radius is halved?

Answer by Boreal(15235)

2pi*r^2+2pi*r
if r=1, then A=2pi+2pi=4pi
if r=3, then A=18pi+6pi=24 pi
6-fold.
---------
A=s^2
A=(3s)^2=9s^2; it increases 9-fold
--------
A=4pi*r^2
A=4pi(1/2*r)^2=p*^r^2, it decreases 75% to 1/4 of the original area.